OpenStudy (anonymous):

A dealer dealt the following cards from a shuffled deck: 3 , 2 , 2 , A , K , Q , K , 5 , 2 , 6 4 , 5 , K , 2 , 7 , 6 , A , J , J , A What was the experimental probability of dealing a black card? can someone help will give medal 8 cards are black and12 are red pleaaaasseee help do i divide 20 by 8 or 8 by 20

OpenStudy (anonymous):

@amistre64 can you help me please

OpenStudy (anonymous):

@mathmate can you help

OpenStudy (anonymous):

@dan815 can you help

OpenStudy (mathmate):

Yes, your answer (8/(12+8)=8/20=2/5) is correct (if there are 8 black and 12 red cards). The experimental probability is the number of observed successes over total number of trials.

OpenStudy (anonymous):

thanks @mathmate could you help with one more

OpenStudy (mathmate):

Just post it and I'll see!

OpenStudy (anonymous):

A businessman bought a car dealership that is incurring a loss of $500,000 a year. He decided to strategize in order to turn the business around. In addition to the $500,000 annual loss, his fixed cost for running the dealership on a monthly basis is $5,000. The number of cars sold per week and their probabilities mimic the outcomes of three coins being flipped. The number of cars sold in a week was observed to be the same as the number of tails that appear when three coins are flipped. See the distribution:

OpenStudy (anonymous):

Given that there are 52 weeks in a year, what is the expected revenue per car (rounded to the nearest dollar) that has to be made in order to break even in the first year

OpenStudy (mathmate):

The distribution (that you have not posted), is it 0 1/8 1 3/8 2 3/8 3 1/8 ?

OpenStudy (anonymous):

yea sorry

OpenStudy (anonymous):

forgot to post that part

OpenStudy (mathmate):

From the distribution, can you find the expectation of the number of tails when we flip three (fair) coins?

OpenStudy (anonymous):

1.25?

OpenStudy (anonymous):

im not good with this:(

OpenStudy (mathmate):

Expectation is defined as: \(E[X]=\sum xP(x)\) Have you learned that before?

OpenStudy (anonymous):

i studied it but it dont make sense

OpenStudy (mathmate):

In this case, E[X]=0(1/8)+1(3/8)+2(3/8)+3(1/8)=0.5

OpenStudy (mathmate):

Sorry, it should read =1.5

OpenStudy (mathmate):

So on the average, the dealer sold 1.5 cars per week.

OpenStudy (mathmate):

or 52*1.5=78 cars per year.

OpenStudy (amistre64):

assume 8 weeks ----------------------- 0 1/8 0 cars are sold 1times out of 8: 0(1) = 0 ----------------------- 1 3/8 1 cars are sold 3times out of 8: 1(3) = 3 ----------------------- 2 3/8 2 cars are sold 3 times out of 8: 2(3) = 6 ------------------------- 3 1/8 3 cars are sold 1 times out of 8: 3(1) = 3 ------------------------- how many cars are sold on average in 8 weeks? 0 + 3 + 6 + 3

OpenStudy (mathmate):

Now what you need to do is to calculate the total loss, including the annual loss, and 12 times the monthly loss. To break even, the dealer has to make profit on each of the 78 cars sold to cover the loss. So divide the total loss by 78 to get the profit needed to break even.

OpenStudy (anonymous):

o+3+6+3 would give the numer of losses

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@mathmate whats the total loss

OpenStudy (mathmate):

" In addition to the $500,000 annual loss, his fixed cost for running the dealership on a monthly basis is $5,000." What would that cost over a year?

OpenStudy (anonymous):

the number of cars sold per week

OpenStudy (mathmate):

..."or 52*1.5=78 cars per year."